OpenStudy (jkbo):

please help 100 is 80% of what number?

OpenStudy (jkbo):

@UnknownRandom are u sure

OpenStudy (unknownrandom):

That seems weird doesn't it?

OpenStudy (jkbo):

yes hahha @UnknownRandom

OpenStudy (unknownrandom):

It's right though. I actually had to check to make sure. Here is a site that might help. http://www.mathsisfun.com/percentage.html

OpenStudy (jkbo):

I mean kit makes sense

OpenStudy (jkbo):

can you hellp me with this one also? Solve the system of equations -7x − 8y = -8 -6x − 9y = -8

OpenStudy (unknownrandom):

Maybe. Give me a min.

OpenStudy (unknownrandom):

I'll let @thu1935 help you.

OpenStudy (anonymous):

call x the number. 100 = 80% of x -> 100 = 80*x/100 -> 8x = 1,000 -> x = 1,000/8 = 125

OpenStudy (jkbo):

@thu1935 thank you I need help with this next problem Solve the system of equations -7x − 8y = -8 -6x − 9y = -8

OpenStudy (unknownrandom):

@thu1935 Thanks for the help. I thought something seemed wrong.

OpenStudy (jkbo):

so what would Y be?

OpenStudy (unknownrandom):

I can't remember how to do any of this right now. @mathslover Can you help with this.

OpenStudy (unknownrandom):

Sorry I went brain dead. Here is a good resource http://www.purplemath.com/modules/systlin5.htm

mathslover (mathslover):

-7x − 8y = -8 --------(1) -6x − 9y = -8 --------(2) In Equation 1 : -7x = 8y - 8 x = \(\cfrac{8y-8}{-7}\) Just put this value in equation 2 and solve for y and then for x

mathslover (mathslover):

It is called substitution method.

OpenStudy (jkbo):

@mathslover where would * go in the second problem

OpenStudy (jkbo):

*8

OpenStudy (anonymous):

-7x - 8y = -8 (times 6) --> -42x - 48y = -48 - 6x - 9y = -8 (times -7) --> 42x + 63y = 56 Add up the 2 equations --> 15y = 8 -> y = 8/15 Replace y by 8/15 in the first equation to get x. -7x - 64/15 = -8 -105x - 64 = -120 --> 105*x = 56 -> x = 56/105

OpenStudy (jkbo):

thank you all

OpenStudy (unknownrandom):

Sorry I couldn't help more. I have been doing too much homework. Have a good night!